Use a grapher to find all local maxima and minima and the values of x where they occur. Give values rounded to the two decimal places.
f(x)=4-x+x^2
g(x)=X^3-4x+1

Answer by ewatrrr(24785)

 
Hi,
Using a 'grapher' to find all local maxima and minima and the values of x where they occur....
Enter the function into the particular graphing Calculator you are Using
and follow directions given for finding maximum and/or minimum points
f(x)=4-x+x^2 Green vertex(1/2, 3.75) minima at 3.75 when x = 1/2
g(x)=X^3-4x+1 Blue